solve the linear system using method of substition.
3m-6n=1
m+3n=2

Answer by mxgirl22(39)

Solve the linear system using method of substitution.
1.) 3m-6n=1
2.) m+3n=2


***NOTE: I numbered each equation so I can refer back to them easier.

--------------------------------------------------
Step 1.) Look at each equation to see which one would be easier to manipulate. It looks like it will be the 2nd one.


1.) 3m-6n=1
2.) m+3n=2

--------------------------------------------------
Step 2.) Get "m" by itself.


2.) m+3n=2
m+3n-3n=2-3n

m=2-3n

--------------------------------------------------
Step 3.) Now that we have "m" by itself in equation 2, we can plug in m=2-3n for the "m" in equation 1.


1.) 3m-6n=1


3(2-3n)-6n=1
6-9n-6n=1
6-15n=1
6-6-15n=1-6
-15n=-5
-15n/-15=-5/-15
n=0.333333(repeating) or n=1/3

--------------------------------------------------
Step 4.) Plug in the value we got for "n" (n=1/3) into equation 2.

2.) m+3n=2

m+3(1/3)=2
m+1=2
m+1-1=2-1
m=1

--------------------------------------------------
Step 5.) Check with original equation(s) 1&2 using the value m=1 and n=1/3.

1.) 3m-6n=1

3(1)-6(1/3)=1
3-6(1/3)=1
3-2=1
1=1 <----- Checks with equation 1.


2.) m+3n=2

1+3(1/3)=2
1+1=2
2=2 <----- Checks with equation 2.


--------------------------------------------------
Step 6.) Answer.

m=1 and n=1/3